How do I write a power as a decimal? Question: 'Write each power as a decimal": Example number one: 2^8. Example number two: 4^5.
Thank you.

Answer by DrBeeee(684)

Just multiply the base by itself n times, where n is the exponent.
(1) 2^8 = 2*2*2*2*2*2*2*2
(2) 2^8 = 256
Or you can use your calculator using the y^x key
Enter 2 in the keyboard, depress the y^x key, enter 8 on the keyboard, then depress enter or equals key. The answer on the display is 256. Since the answer is a whole number a decimal point is not displayed, non-the-less it is the decimal form of the power.
Your second example is
(3) 4^5 = 4*4*4*4*4
(4) 4^5 = 1024
If the questioner insists on the "decimal" your answers are 256. and 1024..